# Evolution of flat bands in MoSe2/WSe2 moiré lattices: A study combining machine learning and band unfolding methods

### v1

@@ III.4 Effect of spin-orbit coupling
We now discuss the effect of SOC on the flat bands. Figure 7 shows the unfolded bands with layer projections onto WSe2 with and without SOC for $H$-6.01 and $R$-6.01. These wavefunctions for the unfolding are obtained from DFT calculations. One can see that the bands near $\Gamma$ remain unchanged with the inclusion of SOC. Therefore, the $\Gamma$-valley flat bands are not affected by SOC. In contrast, there is a SOC splitting of about 0.45 eV at K for both the $H$-6.01 and $R$-6.01 systems. Such a splitting is very close to that for the untwisted MoSe2/WSe2 heterobilayers [Fig. 8(b) in Appendix A]. These features suggest that the SOC splitting is independent of whether it is in the R-stacking or H-stacking for twisting. Also, it has a weak angle dependence. Moreover, one can see that the K-valley band shifts to about 150 meV higher than itself with SOC. Whereas, Fig. 5(f) shows that the $\Gamma$-valley flat band is only 76 meV higher than the K-valley flat band without SOC. Therefore, the K-valley band should be the VBM for MoSe2/WSe2 moiré lattices.
Figure 7: Layer-projected bands with SOC for WSe2 in (a) H-6.01 and (b) R-6.01. The blue and red circles represent the bands with and without SOC, respectively. The band structures are obtained from DFT calculations.
@@ IV CONCLUSIONS
In conclusion, we have investigated the electronic structure of MoSe2/WSe2 moiré lattices by combining machine-learning methods and a band unfolding technique, which enables tracking the evolution of flat bands as a function of the twist angle. We find that multiple flat bands emerge when the twist angle is decreased to 2.13. We reveal that the interlayer coupling and moiré potential are responsible for the emergence of the $\Gamma$-valley flat bands. Whereas, the K-valley flat band is mainly modulated by the long wavelength moiré potentials. Due to this difference, the twist angle for generating the $\Gamma$-valley flat bands is larger than that for the K-valley. And the $\Gamma$-valley flat bands shift to higher energies much faster than the K-valley flat band. By analyzing the characteristics of the wavefunctions of the valence bands at K, we find that the K-valley flat band experiences a triangular-to-hexagonal lattice transition. We have further discussed the effect of SOC on the flat bands of the twisted systems. Our results find that like the untwisted TMD bilayers, the SOC has a minor effect on the $\Gamma$-valley flat bands but induces a significant splitting for the K-valley. By including the SOC, the K-valley flat bands are higher in energy than the $\Gamma$-valley flat bands, that is, the K-valley flat bands contribute to the VBM of the MoSe2/WSe2 moiré lattices.
@@ ACKNOWLEDGMENTS
This work was supported by the National Natural Science Foundation of China (Grants No. 12174098 and No. 11774084).
Appendix A UNTWISTED BILAYER
@@
Appendix B ORBITAL PROJECTIONS
@@
Appendix C R-STACKING MOIRÉ LATTICE
Figures 10(a) and (b) show the geometrical structure of $R$-3.89. The R-1.7 with structural relaxation shows a pattern of triangular networks. Such a pattern is consistent with previous studies and was also seen in the twisted bilayers of other TMD family members  [73, 33, 74, 75, 90, 91, 92, 77, 93, 80, 79, 94, 76, 95]. The displacements of the reconstructed moiré lattice are shown in Figs. 10(c) and (d). Figure 10(e) shows the interlayer distances for the relaxed $R$-1.7 moiré lattice. Figure 10(f) shows the changes in the size of the reconstruction ($\delta_{z}$) and the interlayer distances for high-symmetry sites (RXM, RMX, and RMM) as the twist angle decreases. Overall, they show similar trends with those for H-stacking moiré lattices.
Figures 11(a)-(i) show the evolution of the flat bands of the R-stacking moiré lattice. As the angle decreases from 6.01 to 2.45, the $\Gamma$ valley gradually moves to higher energies than VBMK. A few flat bands emerge when the angle reaches 2.45. Based on the analysis of the flat bands for the H-stacking moiré lattice. One can expect that these bands originate from the $\Gamma$-valley. We further analyzed the wavefunction of the $\Gamma$-valley flat bands of $R$-3.89 [Fig. 11(j)], which supports the above expectation. The wavefunction is mainly localized at RMX and RXM sites with interlayer hybridization. However, the RXM sites have contributed more than RMX. The reason is that unlike the twisted homobilayers the two stackings are no longer symmetrically equivalent.
Figure 11: Band structure of R-stacked moiré lattices. (a)-(i) The valence bands of $R$-$\theta$ ($\theta$ = 6.01, 5.09, 4.41, 3.89, 3.48, 3.15, 2.88, 2.65, 2.45). (j) Charge density distributions of the $\Gamma$ valley (guided by the red diamond) for $R$-3.89 in (d).
There is a difference in the flat bands between H- and R-stacking moiré structures. The bandwidth of the highest $\Gamma$-valley flat band of the R-stacking moiré pattern is larger than that of the H-stacking at the same angle. For example, the bandwidth of the highest $\Gamma$-valley flat band of R-3.89 is 9.4 meV, which is double that for H-3.89 [ 4.6 meV, Fig. 4(d)]. This difference in bandwidth may be related to the localization of high-symmetry stacking configurations. In the H-stacking moiré lattice, the wavefunction of the $\Gamma$-valley flat band is only localized at the HMX site [see Fig. 4(j)]. Whereas the wavefunction of the highest flat band of the R-stacking moiré lattice is localized at the RMX and RXM sites.

### v2

@@ III.4 Effect of spin-orbit coupling
We now discuss the effect of SOC on the flat bands. Figure 10(a) shows the unfolded bands with layer projections onto WSe2 with and without SOC for $H$-6.01 and $R$-6.01. One can see that the bands near $\Gamma$ remain unchanged with the inclusion of SOC. In contrast, there is a SOC splitting of about 0.45 eV at K for both the $H$-6.01 and $R$-6.01 systems. Such a splitting is very close to that for the untwisted MoSe2/WSe2 heterobilayers [Fig. 12(b) in Appendix B]. Note that the $\Gamma$-valley flat band is only 76 meV higher than the K-valley flat band without SOC [Fig. 8(f)]. Therefore, the K-valley band should be the VBM for MoSe2/WSe2 moiré lattices. We further investigate the SOC effect on the band structure of H-3.48, for which the K valley is lower than the $\Gamma$-valley flat band by about 5.8 meV without SOC [Fig. 10(b)]. The inclusion of SOC leads to that the K valley is about 143 meV higher than the $\Gamma$-valley flat band. Moreover, our results also show that the SOC splitting has a weak dependence on twisting and stacking.
Figure 10: Effect of SOC on the band structures of 6.01 and 3.48 moiré lattices. (a) $k$-projected bands with (blue) and without (red) SOC for WSe2 in $H$-6.01 and $R$-6.01. (b) The band structure of $H$-3.48 with band folding. The $\Gamma$-valley flat bands (in red) are placed at the same energy for comparison. The K-valley bands are shown in green.
@@ IV CONCLUSIONS
In conclusion, we have investigated the electronic structure of MoSe2/WSe2 moiré lattices by combining machine-learning methods and a band unfolding technique, which enables tracking the evolution of flat bands as a function of the twist angle. We find that multiple flat bands emerge when the twist angle is decreased to 2.13. We reveal that the interlayer coupling and moiré potential are responsible for the emergence of the $\Gamma$-valley flat bands. Whereas, the K-valley flat band is mainly modulated by the structural reconstruction. Due to this difference, the twist angle for generating the $\Gamma$-valley flat bands is larger than that for the K-valley. And the $\Gamma$-valley flat bands shift to higher energies much faster than the K-valley flat band. By analyzing the characteristics of the wavefunctions of the valence bands at K, we find that the K-valley flat band experiences a triangular-to-hexagonal lattice transition. We have further discussed the effect of SOC on the flat bands of the twisted systems. Our results find that like the untwisted TMD bilayers, the SOC has a minor effect on the $\Gamma$-valley flat bands but induces a significant splitting for the K-valley. By including the SOC, the K-valley flat bands are higher in energy than the $\Gamma$-valley flat bands, that is, the K-valley flat bands contribute to the VBM of the MoSe2/WSe2 moiré lattices.
@@ ACKNOWLEDGMENTS
This work was supported by the National Natural Science Foundation of China (Grants No. 12174098 and No. 11774084) and by the State Key Laboratory of Powder Metallurgy, Central South University, Changsha, China. Calculations were carried out in part using computing resources at the High-Performance Computing Platform of Hunan Normal University.
Appendix A DEEPH MODEL
Figure 11 shows the band structures of $H$-6.01 and $R$-6.01. It should be mentioned that our dataset does not include any moiré structures with the twist angle of 6.01. The band structures obtained by the DeepH model agree well with those from DFT, confirming the high precision of our DeepH model.
Figure 11: Benchmark DeepH calculations of $H$- and $R$-6.01. DFT results are also shown for comparison.
Appendix B UNTWISTED BILAYER
@@
Appendix C ORBITAL PROJECTIONS